# Thermodynamic and observational implications of black holes in toroidal geometry

### v1

@@ VI.2 Null Particles
$r_{\mathrm{ISCO}}$ $r_{\mathrm{ISCO}}$ $r_{\mathrm{ISCO}}$
$Q=1,~\Lambda=-1$ 3.19374 2.63169 3
$Q=1,~\Lambda=0$ $1.50994\times 10^{12}$ 4 3
$Q=1,~\Lambda=1$ 3.20618 0.709427 3
$Q=2,~\Lambda=-1$ 1.46751 1.22795 3
$Q=2,~\Lambda=0$ $1.2595\times 10^{13}$ 2.33181 3
$Q=2,~\Lambda=1$ 1.85118 1.82235 3
We present the behavior of redshift $z_{1}$ (left panel), blueshift $z_{2}$ (middle panel) and gravitation shift $z_{\mathrm{g}}$ (right panel) by setting $M=1,~r_{\mathrm{e}}=r_{\mathrm{ISCO}}$ and compute various curves for different values of $Q$ such as for $Q=0.8$ (red curve), $Q=1$ (blue dotted curve) and $Q=1.3$ (black dotted-dashed curve). It is evident from Fig. 9 that raising the parameter $Q$ leads to higher values of the redshift $z_{1}$ (left panel) and gravitational shift $z_{\mathrm{g}}$ (right panel), while in the case of blueshift $z_{2}$ (middle panel), it decreases with the rise in charge $Q$. It also indicates that as the parameter $Q$ increases, the frequency shift curves grow more rapidly as $\Lambda$ varies. It is worthy to mention here that $z_{\mathrm{g}}$ remain positive and rises with the parameter $Q$ as presented in Fig. 9 (right panel) while on the other hand $z_{2}$ (middle panel) shows negative trend which becomes more negative as the parameter $Q$ rises, owing to the impact of $z_{\mathrm{kin}^{-}}$ component. Thereby, the curves in Fig. 9 represent the highest frequency shifts, computed for an emitter revolving circularly around the BH at $r_{\mathrm{e}}=r_{\mathrm{ISCO}}$, and also as the orbital radius grows infinitely, then the frequency shift vanishes to zero. One can plugged $R=1+z_{1}$ and $B=1+z_{2}$ to verify
$RB$ $=$ $\pi r^{2}_{\mathrm{e}}\left(\frac{3}{6Mr_{\mathrm{e}}-12Q^{2}+\pi\Lambda r^{4}_{\mathrm{e}}}+\frac{2}{8Q^{2}-3Mr_{\mathrm{e}}}\right)~.$ (78)
@@
$M$ $=$ $\frac{\sqrt{r^{2}_{\mathrm{e}}\left(16(RB)^{2}Q^{4}+8\pi(RB)^{2}\Lambda Q^{2}r^{4}_{\mathrm{e}}+\pi^{2}(RB)^{2}\Lambda^{2}r^{8}_{\mathrm{e}}-56\pi RBQ^{2}r^{2}_{\mathrm{e}}-14\pi^{2}RB\Lambda r^{6}_{\mathrm{e}}+\pi^{2}r^{4}_{\mathrm{e}}\right)}}{12RBr^{2}_{\mathrm{e}}}$ (79)
- $\frac{\pi r_{\mathrm{e}}}{12RB}-\frac{\pi\Lambda r^{4}_{\mathrm{e}}-28Q^{2}}{12r_{\mathrm{e}}}\,.$
The highest values of redshift $z_{1}$ and blueshift $z_{2}$ for revolving particles in ISCO are obtained at $r_{\mathrm{ISCO}}$. One can rely on observational data of redshift $z_{1}$ and blueshift $z_{2}$ derived from $\mathrm{H_{2}O}$ megamaser present in the accretion disk surrounding supermassive BHs with active galactic nuclei cores; for example in the case of NGC 4258 glaxy, measurement from its central megamaser system demonstrate that the redshift range lies in $z_{1}\in(4\times 10^{-3},~6\times 10^{-3})$, blueshift range lies in $z_{2}\in(-1.7\times 10^{-3},~-1.0\times 10^{-3})$ and the emitter radius is located in the range of $r_{\mathrm{e}}\in(0.04pc,~0.5pc)$ from the center of the disk Herrnstein:2005xc . The data computed from megamaser glaxies NGC 1194, NGC 2273, NGC 2960, NGC 6264, NGC 6323, and UGC 3789 indicates the redshift range lies in $z_{1}\in(1.1\times 10^{-3},~3.3\times 10^{-3})$, blueshift range is $z_{2}\in(-2.6\times 10^{-3},~-0.9\times 10^{-3})$ and the emitter radius lies in $r_{\mathrm{e}}\in(0.028pc,~1.33pc)$ Villaraos:2022euo ; Reid:2008nm ; Kuo:2010uy . The gravitational redshift in megamaser systems has been shown to range from $z_{\mathrm{g}}\in(1.6\times 10^{-7},~3.1\times 10^{-5})$, governed by the mass of the central supermassive BH and the distance to the emitting region as discussed in Refs. Nucamendi:2020tov ; Villalobos-Ramirez:2022xic ; Villaraos:2022euo ; Gonzalez-Juarez:2022pya . Thereby, one can employ the generic relativistic method to an astrophysical system to compute the gravitational redshift $z_{\mathrm{g}}$.
@@ VII Conclusions
$R^{\mathrm{Rup}}$ Coincide Coincide Coincide both zero points Coincide No-Coincidence
@@
In this analysis, it is notably found that there is only one zero point of heat capacity in the case of the Hawking-Bekenstein and Rènyi, while on the other hand, the exponential corrected entropy demonstrates two distinct zero points of heat capacity. This suggests that exponentially corrected entropy provides a more intricate thermal phase structure and more complex stability behavior. To validate the physical significance of these identified zero points, we employed Weinhold and Ruppeiner geometry, observing that the Ricci scalar obtained from the Ruppeiner geometry diverges at the same point at which heat capacity becomes zero, thereby providing a geometric validation of the phase transition. Compared to the Hawking-Bekenstein and Rènyi entropies, where free energies (HFE and GFE) demonstrate the fluctuating behavior indicating instability, the exponential corrected entropy approach leads to positive and stable behavior in both HFE and GFE, highlighting global thermodynamic stability. Similarly, the exponential corrected entropy model illustrates a non-monotonic pressure trend in the PVP-V diagram, indicating a possible phase transition that is absent in other entropy models, and this is also confirmed by the behavior of isothermal compressibility. Additionally, we presented the impact of entropy corrections on the thermodynamics of charged torus-like BH by comparing it with the spherically symmetric spacetimes such as Schwarzschild and RN-AdS BHs in table 2.
@@
We further studied the sparsity of Hawking radiation and the energy emission rate by utilizing the conjugate temperature, which corresponds to the respective entropy models used in this analysis. It is observed that for small ranges of entropy, the parameter of sparsity $\eta$ increased, but for large values of entropy, it began to decrease. After some time, it became constant for the exponential corrected entropy, but in the case of the Hawking-Bekenstein and Rènyi entropy, it converges to zero. As we mentioned earlier, our aim in this study is to investigate the theoretical (thermodynamics) and observational characteristics of the charged torus-like BH. We analyzed the motion of massive and massless test particles via geodesics to examine the observational aspects of a charged torus-like BH. We analytically determined the redshift, blueshift, and gravitational frequency shift of emitted photons originating from the orbital motion. It is notably observed that, for specific parameter choices such as $Q>1$ and $M\leq 1$, the condition $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$ may not hold for charged torus-like BHs due to their spacetime geometry. We observed that this inequality might hold under the following conditions, such as $M>1$ and $Q\leq 1$. Additionally, our analysis confirms that the calculated values of redshift $z_{1}$ and blueshift $z_{2}$ agree with the observational constraints provided in the megamaser disk system, including NGC 4258 and UGC 3789. The observed consistency suggests that toroidal BH models may be suitable for studying the specific astrophysical BHs, notably when examined through the lens of the frequency shift phenomena. Our analysis highlights the significant role of entropy correction in BH thermodynamics and points out the observational features that arise from the toroidal geometries.
@@
We want to extend this analysis in the future by exploring the rotating toroidal BHs or higher-dimensional geometries to study the role of angular momentum and the impacts of extra dimensions on thermodynamic and observational features. In addition, one can employ different quantum-corrected entropy models, such as Barrow and Sharma-Mittal entropies, to investigate the thermodynamics of BHs, which may offer valuable insights regarding the stability and phase transition in BHs. From an observational perspective, investigating lensing, shadow, and quasinormal modes in torus-like geometries and comparing them with the recent LIGO-Virgo LIGOScientific:2016aoc ; LIGOScientific:2016sjg and EHT data EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xqj may enhance the precision of BH parameter constraints.

### v2

@@ VI.2 Null Particles
$r_{\mathrm{ph}}$      $r_{\mathrm{ISCO}}$ $r_{\mathrm{ph}}$      $r_{\mathrm{ISCO}}$      $r_{\mathrm{ph}}$     $r_{\mathrm{ISCO}}$
Q=0.1,l=0.1Q=0.1,~l=0.1 0.026667   0.178122 2.99332   3.74209 3      6
Q=0.1,l=0.5Q=0.1,~l=0.5 0.026667   0.535088 2.99332   3.74464 3      6
Q=0.1,l=1Q=0.1,~l=1 0.026667   0.853481 2.99332   3.75253 3      6
Q=0.5,l=0.1Q=0.5,~l=0.1 0.666667   0.799668 2.82288   3.53813 3      6
Q=0.5,l=0.5Q=0.5,~l=0.5 0.666667   0.791409 2.82288   3.54078 3      6
Q=0.5,l=1Q=0.5,~l=1 0.666667   0.76026 2.82288   3.54894 3      6
Q=1,l=0.1Q=1,~l=0.1 2.66667   3.19998 1.5    0.225115 3      6
Q=1,l=0.5Q=1,~l=0.5 2.66667   3.19948 1.5    2.59649 3      6
Q=1,l=1Q=1,~l=1 2.66667   3.19792 1.5    2.60651 3      6
Q=1.1,l=0.1Q=1.1,~l=0.1 3.22667   0.264025 1.24206   0.238076 3      6
Q=1.1,l=0.5Q=1.1,~l=0.5 3.22667   0.580468 1.24206   0.488289 3      6
Q=1.1,l=1Q=1.1,~l=1 3.22667   0.810754 1.24206   0.638129 3      6
We present the behavior of redshift $z_{1}$ (upper left panel), blueshift $z_{2}$ (upper right panel) and gravitation shift $z_{\mathrm{g}}$ (lower panel) by setting $M=1,~r_{\mathrm{e}}=r_{\mathrm{ISCO}}$ and compute various curves for different values of $Q$ such as for $Q=0.8$ (red curve), $Q=1$ (blue dotted curve) and $Q=1.3$ (black dotted-dashed curve). It is evident from Fig. 9 that raising the parameter $Q$ leads to higher to lower values of the redshift $z_{1}$ (left panel) (right panel), while in the case of blueshift $z_{2}$ (middle panel), it increases with the lower charge $Q$. It also indicates that, as the parameter $Q$ increases, the frequency-shift curves grow more rapidly with $l$. It is worthy to mention here that $z_{\mathrm{g}}$ increases with neagtive behavior after the AdS radii $l$ increases it becomes positive and rises with the parameter $Q$ as presented in Fig. 9 (lower panel) while on the other hand $z_{2}$ (middle panel) shows negative trend which becomes more negative as the parameter $Q$ rises, owing to the impact of $z_{\mathrm{kin}^{-}}$ component. Furthermore, from the relation of gravitation shift, we observe that it is independent of the term $l$, so we graphically present the behavior of $z_{\mathrm{g}}$ in the form of $r_{\mathrm{e}}$, which indicates that for large values of $r_{\mathrm{e}}$, gravitational shift obtains maximum values. Thereby, the curves in Fig. 9 represent the highest frequency shifts, computed for an emitter revolving circularly around the BH at $r_{\mathrm{e}}=r_{\mathrm{ISCO}}$, and also as the orbital radius grows infinitely, then the frequency shift vanishes to zero. One can plugged $R=1+z_{1}$ and $B=1+z_{2}$ to verify
$RB$ $=$ $\frac{\pi l^{2}r^{2}_{\mathrm{e}}}{l^{2}\left(4Q^{2}-2Mr_{\mathrm{e}}\right)+\pi r^{4}_{\mathrm{e}}}~.$ (78)
@@
$M$ $=$ $\frac{4l^{2}Q^{2}+\pi r^{4}_{\mathrm{e}}}{2l^{2}r_{\mathrm{e}}}-\frac{\pi r_{\mathrm{e}}}{2\text{RB}}\,.$
The highest values of redshift $z_{1}$ and blueshift $z_{2}$ for revolving particles in ISCO are obtained at $r_{\mathrm{ISCO}}$. One can rely on observational data of redshift $z_{1}$ and blueshift $z_{2}$ derived from $\mathrm{H_{2}O}$ megamaser present in the accretion disk surrounding supermassive BHs with active galactic nuclei cores; for example in the case of NGC 4258, measurement from its central megamaser system demonstrate that the redshift range lies in $z_{1}\in(4\times 10^{-3},~6\times 10^{-3})$, blueshift range lies in $z_{2}\in(-1.7\times 10^{-3},~-1.0\times 10^{-3})$ and the emitter radius is located in the range of $r_{\mathrm{e}}\in(0.04pc,~0.5pc)$ from the center of the disk Herrnstein:2005xc . The data computed from megamaser glaxies NGC 1194, NGC 2273, NGC 2960, NGC 6264, NGC 6323, and UGC 3789 indicates the redshift range lies in $z_{1}\in(1.1\times 10^{-3},~3.3\times 10^{-3})$, blueshift range is $z_{2}\in(-2.6\times 10^{-3},~-0.9\times 10^{-3})$ and the emitter radius lies in $r_{\mathrm{e}}\in(0.028pc,~1.33pc)$ Villaraos:2022euo ; Reid:2008nm ; Kuo:2010uy . The gravitational redshift in megamaser systems has been shown to range from $z_{\mathrm{g}}\in(1.6\times 10^{-7},~3.1\times 10^{-5})$, governed by the mass of the central supermassive BH and the distance to the emitting region as discussed in Refs. Nucamendi:2020tov ; Villalobos-Ramirez:2022xic ; Villaraos:2022euo ; Gonzalez-Juarez:2022pya . Thereby, one can employ the generic relativistic method to an astrophysical system to compute the gravitational redshift $z_{\mathrm{g}}$.
@@ VII Conclusions
$R^{\mathrm{Rup}}$ Coincide Coincide Coincide both ZPs Coincide No-Coincidence
@@
In this analysis, it is notably found that there is only one ZP of heat capacity in the case of the Hawking-Bekenstein and Rènyi, while on the other hand, the exponential corrected entropy demonstrates two distinct ZPs of heat capacity. This suggests that exponentially corrected entropy provides a more intricate thermal phase structure and more complex stability behavior. To validate the physical significance of these identified ZPs, we employed Weinhold and Ruppeiner geometries, observing that the Ricci scalar obtained from the Ruppeiner geometry diverges at the same point where heat capacity becomes zero, thereby providing geometric validation of the phase transition. Compared to the Hawking-Bekenstein and Rènyi entropies, where free energies (HFE and GFE) demonstrate the fluctuating behavior indicating instability, the exponential corrected entropy approach leads to positive and stable behavior in both HFE and GFE, highlighting global thermodynamic stability. Similarly, the exponential corrected entropy model shows a non-monotonic pressure trend in the PVP-V diagram, indicating a possible phase transition absent in other entropy models, and this is also confirmed by the behavior of isothermal compressibility. Additionally, we presented the impact of entropy corrections on the thermodynamics of charged torus-like BH by comparing it with the spherically symmetric spacetimes such as Schwarzschild and RN-AdS BHs in table 2.
@@
We further studied the sparsity of Hawking radiation and the energy emission rate by utilizing the conjugate temperature, which corresponds to the respective entropy models used in this analysis. It is observed that for small ranges of entropy, the parameter of sparsity $\eta$ increased, but for large values of entropy, it began to decrease. After some time, it became constant for the exponential corrected entropy, but in the case of the Hawking-Bekenstein and Rènyi entropy, it converges to zero. As we mentioned earlier, our aim in this study is to investigate the theoretical (thermodynamics) and observational characteristics of the charged torus-like BH. We analyzed the motion of massive and massless test particles via geodesics to examine the observational aspects of a charged torus-like BH. We analytically determined the redshift, blueshift, and gravitational frequency shift of emitted photons originating from the orbital motion. It is notably observed that, for specific parameter choices such as $Q>1$ and $M\leq 1$, the condition $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$ may not hold for charged torus-like BHs due to their spacetime geometry. We observed that this inequality might hold under the following conditions, such as $M>1$ and $Q\leq 1$. Additionally, our analysis confirms that the calculated redshift $z_{1}$ and blueshift $z_{2}$ agree with the observational constraints for the megamaser disk systems NGC 4258 and UGC 3789. The observed consistency suggests that toroidal BH models may be suitable for studying specific astrophysical BHs, notably when examined through the lens of frequency-shift phenomena. Our analysis highlights the significant role of entropy correction in BH thermodynamics and identifies observational features arising from toroidal geometries.
@@
We want to extend this analysis in the future by exploring rotating toroidal BHs and higher-dimensional geometries to study the role of angular momentum and the impact of extra dimensions on thermodynamic and observational features. In addition, one can employ different quantum-corrected entropy models, such as the Barrow and Sharma-Mittal entropies, to investigate the thermodynamics of BHs, which may offer valuable insights into their stability and phase transitions. From an observational perspective, investigating lensing, shadow, and quasinormal modes in torus-like geometries and comparing them with the recent LIGO-Virgo LIGOScientific:2016aoc ; LIGOScientific:2016sjg and EHT data EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xqj may enhance the precision of BH parameter constraints.